OpenStudy (anonymous):

Solve by factoring: 5x2 = –45x

OpenStudy (anonymous):

5x2 = –45x 5x2 +45x = 0 5x(x+9)=0

hero (hero):

5x^2 = -45x x^2 = -9x x^2 + 9x = 0 x(x + 9) = 0 x = 0, -9

OpenStudy (anonymous):

THank you guys!!